# cuPSS: a package for pseudo-spectral integration of stochastic PDEs

### v1

@@ Abstract
A large part of modern research, especially in the broad field of complex systems, relies on the numerical integration of PDEs, with and without stochastic noise. This is usually done with eiher in-house made codes or external packages like MATLAB, Mathematica, Fenicsx, OpenFOAM, Dedalus, and others. These packages rarely offer a good combination of speed, generality, and the option to easily add stochasticity to the system, while in-house codes depend on certain expertise to obtain good performance, and are usually written for each specific use case, sacrificing modularity and reusability. This paper introduces a package written in CUDA C++, thus enabling by default gpu acceleration, that performs pseudo-spectral integration of generic stochastic PDEs in flat lattices in one, two and three dimensions. This manuscript describes how the basic functionality of cuPSS, with an example and benchmarking, showing that cuPSS offers a considerable improvement in speed over other popular finite-difference and spectral solvers.
preprint: APS/123-QED
@@ I.1 Pseudo-spectral integration
Computing the right-hand side numerically in a system discretized in a lattice of size $N$ would thus require $O(N^{3})$ calculations; $N^{2}$ calculations for the convolution at each wavevector $\mathbf{q}$. A pseudo-spectral method consists of viewing the previous equation in the following equivalent way
@@ I.2 De-aliasing
One last detail must be mentioned when talking about these numerical methods, for which we need to introduce an explicit discretization of space. Consider a field $\phi(x,t)$, defined in a one-dimensional space of size $N$, discretized in $N/\Delta x$ points, each separated from the next at a distance of $\Delta x$. This induces a natural discretization of Fourier space in $N$ frequencies $q_{i}$ at a distance of $\Delta q=2\pi/(N\Delta x)$, where it is common to choose a frequency interval symmetric around $q=0$, such that $q_{i}\in(-\pi/\Delta x,\pi/\Delta x)$. Let us assume for simplicity, and without loss of generality, that $\Delta x=1$. We then have a lattice of $N$ frequencies going from $-\pi$ to $\pi$. Let us now consider the discrete Fourier series of the field $\phi(x)$, which can be written $\phi(x)=\sum_{q_{i}}\tilde{\phi}(q_{i})\exp(iq_{i}x)$. Consider a quadratic nonlinearity derived from this field
@@ V Callback functions
cuPSS offers direct dynamic access to all fields and terms, meaning the user can access directly the value of all fields during the progress of a simulation. It offers an easy way to do so through callback functions; functions that are called giving access to the real values of a field at the end of a timestep. There are many applications for callback functions, such as applying boundary conditions. Since a Fourier spectral decomposition assumes periodic boundary conditions, applying other boundary conditions through callback functions should be done with caution. Setting Dirichlet boundary conditions, which would be properly done spectrally through the choice of a different spectral basis [25], can then be done, for instance, by setting a field to a certain value at a range of positions close to the boundary, making sure this range is bigger than all lengthscales of the problem.
@@ VI Benchmarking
This section shows limited benchmarking of cuPSS in different systems, a consumer level system with a Intel Core i7-13700K CPU and a NVIDIA RTX 3060 Ti GPU, and a high-performance computing cluster, running in a few different GPUs. For reference, some benchmarks are compares to the same systems integrated in two different packages, FEniCSx [29], a finite difference solver for PDEs written in weak form, and Dedalus [30], another spectral solver that implements the same spectral techniques as cuPSS, together with extra features such as non-flat spaces and other spectral basis. This section shows the main advantage of cuPSS, which achieves much higher speeds compared to other packages through the direct implementation of the library on CUDA C++, and having a simplified codebase that relies more on speed than depth of features and breadth.
The benchmarks have been run with two example solvers, for models B and H [3], as these two solvers implement both a simple example with a single field and nonlinear term, and a more complex example with several fields and constraints. The benchmark results are given in terms of average time taken to integrate one timestep, calculated as an average over a number of timesteps that ranges from 100 to 10000 depending on the problem and system size. If there is any overhead in any of these solvers, associated to preparing the problem, reserving memory and so on, this time has not been counted.
@@
Figure 4: Same as Fig. 3, but for a solver for model H, considerably more complicated and with several intermediate fields to calculate. In the case of cuPSS, the pressure and explicit refer to solvers that solve the Poisson equation for the pressure and the flow directly, as described in the Examples section.
@@
There is one series of benchmarking done on CPU, while all others run on GPUs. CPUs are only more efficient when integrating small systems, in this case when the system size is less than $128\times 128$. This corresponds to 16384 lattice points, which means that one dimensional systems smaller than this should typically run faster on CPU. Due to the limited extent of this benchmarking, every solver should be tested on both to verify which option offers a higher speed. Benchmarks have also been run only on system sizes that are powers of 2. These system sizes result in the best speeds given the nature of both the way CUDA launches kernels on GPUs, and the way FFT algorithms work. This means running on system sizes different than powers of 2 might show a loss of speed compared to the system size. In the worst case scenario, a system size different than $2^{n}$ might run at the same speed as a system size that is equal to the next power of $2$, specially when running on GPUs.
@@ VII Future development
With continuum development and optimization, cuPSS has the potential to become a useful tool for researchers in very different areas, as it is already being used in work, specially in its purely deterministic form, some of which has already been published [32]. Written in the CUDA C++ standard, it should be accessible to most research groups to, not only use, but easily modify and adapt to their needs.
@@ Acknowledgements.
I thank Paarth Gulati for help testing during the early parts of the development of cuPSS, and for suggesting its name, and Daniel Hellstein for guidance in the development of the equation parser. This work was done during stays at the University of California, Santa Barbara and Brandeis University, and was partially funded by NSF grants DMR-2041459 & DMR-2011846.

### v2

@@ Abstract
A large part of modern research, especially in the broad field of complex systems, relies on the numerical integration of PDEs, with and without stochastic noise. This is usually done with eiher in-house made codes or external packages like MATLAB, Mathematica, Fenicsx, OpenFOAM, Dedalus, and others. These packages rarely offer a good combination of speed, generality, and the option to easily add stochasticity to the system, while in-house codes depend on certain expertise to obtain good performance, and are usually written for each specific use case, sacrificing modularity and reusability. This paper introduces a package written in CUDA C++, thus enabling by default gpu acceleration, that performs pseudo-spectral integration of generic stochastic PDEs in flat lattices in one, two and three dimensions. This manuscript describes the basic functionality of cuPSS, with an example and benchmarking, showing that cuPSS offers a considerable improvement in speed over other popular finite-difference and spectral solvers.
@@ I.1 Pseudo-spectral integration
Computing the right-hand side numerically in a system discretized in a lattice of size $N$ would thus require $O(N^{3})$ calculations; $O(N^{2})$ calculations for the convolution at each wavevector $\mathbf{q}$. A pseudo-spectral method consists of viewing the previous equation in the following equivalent way
@@ I.2 De-aliasing
One last detail must be mentioned when talking about these numerical methods, for which we need to introduce an explicit discretization of space. Consider a field $\phi(x,t)$, defined in a one-dimensional space of size $N$, discretized in $N/\Delta x$ points, each separated from the next at a distance of $\Delta x$. This induces a natural discretization of Fourier space in $N$ frequencies $q_{i}$ with a spacing of $\Delta q=2\pi/(N\Delta x)$, where it is common to choose a frequency interval symmetric around $q=0$, such that $q_{i}\in(-\pi/\Delta x,\pi/\Delta x)$. Let us assume for simplicity, and without loss of generality, that $\Delta x=1$. We then have a lattice of $N$ frequencies going from $-\pi$ to $\pi$. Let us now consider the discrete Fourier series of the field $\phi(x)$, which can be written $\phi(x)=\sum_{q_{i}}\tilde{\phi}(q_{i})\exp(iq_{i}x)$. Consider a quadratic nonlinearity derived from this field
@@ V Callback functions
cuPSS offers direct dynamic access to all fields and terms, meaning the user can access directly the value of all fields during the progress of a simulation. It offers an easy way to do so through callback functions; functions that are called giving access to the real values of a field at the end of a timestep. There are many applications for callback functions, such as applying boundary conditions. Since a Fourier spectral decomposition assumes periodic boundary conditions, applying other boundary conditions through callback functions should be done with caution. Setting Dirichlet boundary conditions, which would be properly done spectrally through the choice of a different spectral basis [25], is also usually done by setting a buffer zone of infinite stiffness in the boundaries of the lattice [29, 30], i.e. by setting a field to a certain value at a range of positions close to the boundary, making sure this range is bigger than all lengthscales of the problem.
@@ VI Benchmarking
This section shows limited benchmarking of cuPSS in different systems, a consumer level system with a Intel Core i7-13700K CPU and a NVIDIA RTX 3060 Ti GPU, and a high-performance computing cluster, running in a few different GPUs. For reference, some benchmarks are compared to the same systems integrated in two different packages, FEniCSx [31], a finite difference solver for PDEs written in weak form, and Dedalus [32], another spectral solver that implements the same spectral techniques as cuPSS, together with extra features such as non-flat spaces and other spectral basis. This section shows the main advantage of cuPSS, which achieves much higher speeds compared to other packages through the direct implementation of the library on CUDA C++, and having a simplified codebase that relies more on speed than depth of features and breadth.
The benchmarks have been run with two example solvers, for models B and H [3], as these two solvers implement both a simple example with a single field and nonlinear term, and a more complex example with several fields and constraints. The model B solver is as described in the Example section, while the model H solver is supplied in the package’s repository. The benchmark results are given in terms of average time taken to integrate one timestep, calculated as an average over a number of timesteps that ranges from 100 to 10000 depending on the problem and system size. If there is any overhead in any of these solvers, associated to preparing the problem, reserving memory and so on, this time has not been taken into account in the time measurements.
@@
Figure 4: Same as Fig. 3, but for a solver for model H, considerably more complicated and with several intermediate fields to calculate. In the case of cuPSS, the pressure and explicit refer to solvers that solve the Poisson equation for the pressure and the flow directly.
@@
There is one series of benchmarking done on CPU, while all others run on GPUs. CPUs are only more efficient when integrating small systems, in this case when the system size is less than $128\times 128$. Due to the limited extent of this benchmarking, it is good practive to test any particular solver on different CPUs and GPUs to verify which option offers a higher speed. Benchmarks have also been run only on system sizes that are powers of 2. These system sizes result in the best speeds given the nature of both the way CUDA launches kernels on GPUs, and the way FFT algorithms work. This means running on system sizes different than powers of 2 might show a loss of speed compared to the system size. In the worst case scenario, a system size different than $2^{n}$ might run at the same speed as a system size that is equal to the next power of $2$, specially when running on GPUs.
@@ VII Future development
With continuum development and optimization, cuPSS has the potential to become a useful tool for researchers in very different areas, as it is already being used in work, specially in its purely deterministic form, some of which has already been published [34, 35, 29]. Written in the CUDA C++ standard, it should be accessible to most research groups to, not only use, but easily modify and adapt to their needs.
@@ Acknowledgements.
I thank Paarth Gulati for help testing during the early parts of the development of cuPSS, and for suggesting its name, and Daniel Hellstein for guidance in the development of some of the solver components. This work was done during stays at the University of California, Santa Barbara and Brandeis University, and was partially funded by NSF grants DMR-2041459 & DMR-2011846. This work was also supported by the NSF through DMR 2309635 and the Brandeis Center for Bioinspired Soft Materials, an NSF MRSEC (DMR-2011846). We also acknowledge computational support from NSF XSEDE computing resources allocation TG-MCB090163, National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award BES-ERCAP0026774, and the Brandeis HPCC which is partially supported by the NSF through DMR-MRSEC 2011846 and OAC-1920147.